# Taxonomy-Guided Zero-Shot Recommendations with LLMs

### v2

@@ Abstract
With the emergence of large language models (LLMs) and their ability to perform a variety of tasks, their application in recommender systems (RecSys) has shown promise. However, we are facing significant challenges when deploying LLMs into RecSys, such as limited prompt length, unstructured item information, and un-constrained generation of recommendations, leading to sub-optimal performance. To address these issues, we propose a novel method using a taxonomy dictionary. This method provides a systematic framework for categorizing and organizing items, improving the clarity and structure of item information. By incorporating the taxonomy dictionary into LLM prompts, we achieve efficient token utilization and controlled feature generation, leading to more accurate and contextually relevant recommendations. Our Taxonomy-guided Recommendation (TaxRec) approach features a two-step process: one-time taxonomy categorization and LLM-based recommendation, enabling zero-shot recommendations without the need for domain-specific fine-tuning. Experimental results demonstrate TaxRec significantly enhances recommendation quality compared to traditional zero-shot approaches, showcasing its efficacy as personal recommender with LLMs. Code is available at https://github.com/yueqingliang1/TaxRec.
@@
Yueqing Liang1, Liangwei Yang2, Chen Wang2, Xiongxiao Xu1, Philip S. Yu2, Kai Shu1 1Illinois Institute of Technology, Chicago, USA 2University of Illinois at Chicago, Chicago, USA {yliang40, xxu85}@hawk.iit.edu {lyang84, cwang266, psyu}@uic.edu kshu@iit.edu
@@ 1 Introduction
Due to the emergent ability Wei et al. (2022), large language models (LLMs) have triggered the purse of artificial general intelligence Fei et al. (2022), where an artificial intelligence (AI) system can solve numerous tasks. Tasks that were previously completed separately are now combined into one language modeling task by using prompt templates to turn them into sentences. As shown in Figure 1(a), one single LLM Achiam et al. (2023) can act as our personal assistant to complete a series of tasks such as question answering Tan et al. (2023), machine translation Zhang et al. (2023a) and grammar checking Yasunaga et al. (2021). Besides, LLM-based assistant can also provide reasonable recommendations with its own knowledge within the pre-trained parameters Gao et al. (2023). Without the need for fine-tuning on historical user-item interactions, it acts as the zero-shot recommenders, which greatly extends LLMs toward a more generalized all-task-in-one AI assistant.
@@
Figure 1: (a) LLMs are zero-shot recommenders without knowing other user-item interactions when acting as personal recommendation assistant. (b) Three challenges occur when integrating RecSys within the all-task-in-one LLM assistant, i.e., limited prompt length, vague item title and unconstrained generation.
@@
Acting as the assistant for recommendation, LLMs face several challenges when it meets the requirement from recommender system (RecSys) as shown in Figure 1(b). (1) Limited prompt length prohibits input of all items. In RecSys, the size of item pool effortlessly grows over millions with each represented by tens of tokens, which easily surpasses the prompt length limit Pal et al. (2023) of LLMs. Let alone the long context also causes decoding problems Liu et al. (2024) even the whole item pool is small enough to fit within the prompt. (2) Vague and unstructured item title and description. The text information of items is provided at the will of merchant, which is usually unstructured and vague Ni et al. (2019) to understand without sufficient contexts. As shown in Figure 1(b), the title "1984" can represent the year/book/movie and "Emma" is able to represent people name/book. Direct recommendation with the raw item titles can suffer from the ambiguity prompt issue and leads to inferior performance. (3) Un-constrained generation out of candidate item pools. The generation process of LLMs is un-constrained, and can easily be un-matchable within the item pool, especially for the unstructured titles. For example, the LLMs can generate an item "Punch-Out!!!" that totally out of the item pool when we only provide user’s historical interactions. With the direct text-based generation, it is also compute intensive and mostly infeasible to calculate the ranking score for all candidate items within the pool.
@@
In this paper, we propose using taxonomy dictionary to cope with previously mentioned challenges. Taxonomy is able to provide a systematic framework for identifying, organizing, and grouping items. For each dataset, we first retrieve the taxonomy dictionary from LLM to obtain the categorization knowledge of the domain and categorize all candidate items into structured item pool. It alleviates the vague item title/description problem by providing more item context information. For example, the item "1984" is more clarified as "Type: Book, Genre: Fiction, Theme: Power, …". When prompting LLM for recommendation, we add the taxonomy dictionary within the prompt to inform LLM the candidate items information.
@@
The taxonomy dictionary is a condensed categorization of whole item pool. Compared with adding all candidate items, adding the dictionary can greatly save the tokens needed to inform LLM the candidates information, alleviating the limited prompt length challenge. Instead of directly generating tokens within the item title, we propose to generate categorized features from the taxonomy dictionary. As the taxonomy dictionary can be easily feed within the prompt, it is more controllable to generate features within the dictionary with our designed prompt template. We finally calculate the feature matching score within the categorized item pool to rank the items for recommendation.
@@
Our taxonomy-based approach is a two step process. The first is a one-time taxonomy categorization step, which retrieves knowledge from LLM to build taxonomy and categorized item pool. The second is LLM-based Recommendation step to inference user’s preference based on the historical interactions. This approach effectively handles large item pools, making it feasible to work within LLM token limits, leading to a more efficient, accurate, and scalable recommendation process. Our contributions are summarized as:
@@
Figure 2: The proposed TaxRec for zero-shot LLM-based recommendation. (a) One-time Taxonomy Categorization step aims to generate in-domain taxonomy and enrich/categorize item’s title into structured text information. (b) LLM-based Recommendation step provide ranked item lists for users based on the user’s historical interactions.
@@ 2.1 LLM for Recommendations
Recommendation systems play a critical role in assisting users in finding relevant and personalized items or content. With the emergence of LLMs in recent years, there has been a growing interest in harnessing the power of these models to enhance recommendation systems. LLM-based recommender systems can be mainly categorized into two types: discriminative and generative Wu et al. (2023). The core idea of discriminative LLM-based recommender systems is to utilize discriminative LLMs to learn better representations of users and items from contextual information for downstream tasks Hou et al. (2022); Li et al. (2023); Xiao et al. (2022); Zhang et al. (2022); Yuan et al. (2023); Yao et al. (2022). Unlike discriminative methods, generative LLM-based recommendation systems leverage the generative ability of LLMs to make recommendations. They achieve this by translating the traditional ranking-based recommendation tasks into natural language tasks. Compared with score computing and ranking strategies, they apply techniques such as prompt tuning and in-context learning to make LLMs able to generate the recommendations directly. Geng et al. (2022) is one of the first text-to-text generative recommendation methods, which used the pre-trained T5 as the backbone. Recently, many methods explored the possibility of using LLMs to generate recommendations without fine-tuning Dai et al. (2023); Gao et al. (2023); Liu et al. (2023); Lyu et al. (2023); Wang et al. (2023c). Wang et al. (2023b) integrated database to LLMs to serve as autonomous agents for multiple recommendation tasks. Wang et al. (2023a) used LLMs to augment the user-item interaction graph for recommendation. Lyu et al. (2023) conducted a comprehensive investigation on how different types of prompts will influence the recommendation results. While Wang et al. (2023c) proposed a multi-round self-reflection framework for LLM-based sequential recommendation. However, previous works did not consider the problem that text representations of items are vague and unstructured, which will hinder LLM’s recommendation ability. In this paper, we propose a taxonomy-guided framework to solve this problem.
@@ 2.2 Zero-shot Recommendations
Zero-shot recommendation has emerged as a significant area in the field of recommendation systems. Unlike traditional systems that rely on model training, zero-shot recommendation systems aim to predict user preferences without parameter adjustment. Various methodologies have been explored to address this challenge. For instance, a content-based approach leverages item attributes Lian et al. (2018); Zhang et al. (2023b). Other methods utilize pre-trained language models and embeddings to capture item and user characteristics from textual information Ding et al. (2022); Hou et al. (2022); Li et al. (2023). With the rise of LLMs in recent years, several studies have explored their potential for making zero-shot recommendations Hou et al. (2024); Wang and Lim (2023); He et al. (2023); Feng et al. (2024); Wang et al. (2023b). However, these studies faced the issue of limited context length in LLMs, which prevents the input of all items. Some approaches addressed this by using external tools or information Wang et al. (2023b); Feng et al. (2024), while others employed a plug-in recommendation model to pre-select a small candidate pool Hou et al. (2024); Wang and Lim (2023); He et al. (2023). None of these methods have been able to solve this problem solely using the knowledge from LLMs. This paper proposes a taxonomy-guided LLM recommender that compresses the entire item pool using an LLM-generated taxonomy, enabling zero-shot recommendations without relying on external knowledge.
@@ 3.1 Problem Formulation
Without other user-item interactions, LLMs act as zero-shot recommenders when users directly seek recommendations. The task is to generate the Tok-k recommended items $i$s from the candidate item pool =$\mathcal{I}=\{i_{j}\}_{j=1}^{|\mathcal{I}|}$ only based on user’s historical interactions =$\mathcal{H}=\{i_{1},i_{2},...,i_{|\mathcal{H}|}\}$ and the knowledge within LLMs. As a pure text-based approach, each item $i$ is a title string as shown in Figure 2. The task can be then represented as designing a LLM-based function as:
@@ 3.2 One-time Taxonomy Categorization
The first step is a one-time generation, which aims to structure and clarify items into a categorized item pool. The original item text representation is vague and unstructured, which poses challenges for LLMs to understand and infer user’s interest. As the first item within the pool shown in Figure 2(a), "1984" can be represented as either year/book/movie. Without sufficient in-domain background knowledge, direct recommendation in zero-shot manner with these vague and unstructured textual information is challenging for LLMs.
@@
where $P_{\text{Categorization}}$ is the Categorization Prompt as shown in Table 1 to obtain $i$’s categorized feature list as $i^{C}=[f_{1},f_{2},...,f_{|i^{C}|}]$. We can structure and enrich item textual descriptions with knowledge from LLMs. For example, as shown in the categorized item pool in Figure 2(a), the book "1984" is enriched with "fiction" as genre and "power" as theme. Compared with the origional vague book title, the enriched texts provide more detailed information to assist LLMs inference user’s interests. The categorized item pool $\mathcal{I}^{C}$ is obtained by categorizing items in $\mathcal{I}$ with Equation 3.
@@ 3.3 LLM-based Recommendation
where $s$ is the categorized recommendation, which is a text sequence of key-value pairs representing item’s features within $\mathcal{T}$. $P_{\text{Recommendation}}$ is the recommendation prompt given $\mathcal{H^{C}}$ and $\mathcal{T}$as shown in Table 1. Using $\mathcal{T}$ instead of the item pool can greatly decrease the prompt length and fits the in-domain item’s information within the limited context requirement from LLMs. $P_{\text{Recommendation}}$ also regularizes LLM’s generation format as a list of features based on $\mathcal{T}$. $s$ is further parsed as the feature list $F=[f_{1},f_{2},...,f_{|F|}]$ representing recommended features. Then the ranking score of each item $i$ is calculated as:
@@
Then items with Top-k highest ranking scores are retrieved from item pool and recommended to users. In summary, we designed a framework TaxRec, which uses a taxonomy as the intermediate, to unify the representation of items throughout the recommendation pipeline. TaxRec can retrieve LLM’s knowledge for zero-shot recommendations without any training and other users’ interactions with item.
@@ 4.1.1 Datasets
Table 2: Performance comparison between different zero-shot recommendation baselines and TaxRec. We report Recall and NDCG @(1, 5, 10) results multiplied by 10. The boldface indicates the best result and the underlined indicates the second best. All TaxRec results are significantly better than the baselines with t-test $p<0.05$.
Datasets Methods Recall@1 Recall@5 Recall@10 NDCG@1 NDCG@5 NDCG@10 Movie Popularity 0.005 0.035 0.160 0.005 0.020 0.061 AvgEmb 0.000 0.040 0.100 0.000 0.020 0.039 ZESRec 0.032 0.095 0.222 0.032 0.059 0.099 UniSRec 0.032 0.063 0.143 0.032 0.048 0.074 RecFormer 0.016 0.141 0.219 0.016 0.077 0.103 DirectRec 0.045 0.100 0.180 0.045 0.074 0.099 TaxRec 0.060 0.175 0.300 0.060 0.117 0.157 Book Popularity 0.030 0.070 0.155 0.030 0.046 0.073 AvgEmb 0.005 0.075 0.115 0.005 0.038 0.051 ZESRec 0.005 0.070 0.115 0.005 0.037 0.051 UniSRec 0.000 0.050 0.085 0.000 0.025 0.035 RecFormer 0.010 0.060 0.125 0.010 0.033 0.054 DirectRec 0.000 0.015 0.025 0.000 0.006 0.010 TaxRec 0.070 0.150 0.240 0.070 0.109 0.138

### v3

@@ Abstract
With the emergence of large language models (LLMs) and their ability to perform a variety of tasks, their application in recommender systems (RecSys) has shown promise. However, we are facing significant challenges when deploying LLMs into RecSys, such as limited prompt length, unstructured item information, and un-constrained generation of recommendations, leading to sub-optimal performance. To address these issues, we propose a novel Taxonomy-guided Recommendation (TaxRec) framework to empower LLM with category information in a systematic approach. Specifically, TaxRec features a two-step process: one-time taxonomy categorization and LLM-based recommendation. In the one-time taxonomy categorization phase, we organize and categorize items, ensuring clarity and structure of item information. In the LLM-based recommendation phase, we feed the structured items into LLM prompts, achieving efficient token utilization and controlled feature generation. This enables more accurate, contextually relevant, and zero-shot recommendations without the need for domain-specific fine-tuning. Experimental results demonstrate that TaxRec significantly enhances recommendation quality compared to traditional zero-shot approaches, showcasing its efficacy as a personal recommender with LLMs. Code is available at: https://github.com/yueqingliang1/TaxRec.
@@
Yueqing Liang1, Liangwei Yang2, Chen Wang2, Xiongxiao Xu1, Philip S. Yu2, Kai Shu3thanks: Corresponding author 1Illinois Institute of Technology, Chicago, IL 2University of Illinois at Chicago, Chicago, IL 3Emory University, Atlanta, GA {yliang40, xxu85}@hawk.iit.edu {lyang84, cwang266, psyu}@uic.edu kai.shu@emory.edu
@@ 1 Introduction
Due to the emergent ability Wei et al. (2022), large language models (LLMs) have triggered the purse of artificial general intelligence (AGI) Fei et al. (2022), where an artificial intelligence (AI) system can solve numerous tasks. Tasks that were previously completed separately are now combined into one language modeling task by using prompt templates to turn them into sentences. As shown in Figure 1(a), one single LLM Achiam et al. (2023) can act as our personal assistant to complete a series of tasks such as question answering Tan et al. (2023), machine translation Zhang et al. (2023a) and grammar checking Yasunaga et al. (2021). Besides, an LLM-based assistant can also provide reasonable recommendations with its own knowledge within the pre-trained parameters Gao et al. (2023). Without the need for fine-tuning on historical user-item interactions, it acts as the zero-shot recommenders, which greatly extends LLMs toward a more generalized all-task-in-one AI assistant.
@@
Figure 1: (a) LLMs are zero-shot recommenders without knowing other user-item interactions when acting as personal recommendation assistants. (b) Three challenges occur when integrating RecSys within the all-task-in-one LLM assistant, i.e., limited prompt length, vague item title and unconstrained generation.
@@
Acting as the assistant for recommendation, LLMs face several challenges when it meets the requirement from recommender system (RecSys) as shown in Figure 1(b). (1) Limited prompt length prohibits input of all items. In RecSys, the size of item pool effortlessly grows over millions with each represented by tens of tokens, which easily surpasses the prompt length limit Pal et al. (2023) of LLMs. Let alone the long context also causes decoding problems Liu et al. (2024) even the whole item pool is small enough to fit within the prompt. (2) Vague and unstructured item title and description. The text information of items is provided at the will of merchant, which is usually unstructured and vague Ni et al. (2019) to understand without sufficient contexts. As shown in Figure 1(b), the title “1984” can represent the year/book/movie and “Emma” is able to represent people name/book. Direct recommendation with the raw item titles can suffer from the ambiguity prompt issue and leads to inferior performance. (3) Un-constrained generation out of candidate item pools. The generation process of LLMs is un-constrained, and can easily be un-matchable within the item pool, especially for the unstructured titles. For example, the LLMs can generate an item “Punch-Out!!!” that totally out of the item pool when we only provide user’s historical interactions. With the direct text-based generation, it is also compute intensive and mostly infeasible to calculate the ranking score for all candidate items within the pool.
@@
In this paper, we propose leveraging a taxonomy dictionary to address the aforementioned challenges. A taxonomy provides a systematic framework for identifying, organizing, and grouping items effectively. For each dataset, we first retrieve an LLM to obtain a taxonomy dictionary that contains the categorization knowledge of a domain. This then enables us to categorize all candidate items into a structured item pool, thereby mitigating the issue of vague item titles or descriptions by providing richer item contextual information. For instance, an item like “1984” can be clarified as “Type: Book, Genre: Fiction, Theme: Power, …”. When prompting the LLM for recommendations, we incorporate the taxonomy dictionary into the prompt to enrich the model’s understanding of the candidate items and their attributes.
@@
The taxonomy dictionary is a condensed categorization of the whole item pool. Compared with adding all candidate items, adding the dictionary can greatly save the tokens needed to inform LLM the candidates information, alleviating the limited prompt length challenge. Instead of directly generating tokens within the item title, we propose to generate categorized features from the taxonomy dictionary. As the taxonomy dictionary can be easily fed within the prompt, it is more controllable to generate features within the dictionary with our designed prompt template. We finally calculate the feature matching score within the categorized item pool to rank the items for recommendation.
@@
Our taxonomy-based approach is a two-step process. The first is a one-time taxonomy categorization step, which retrieves knowledge from LLM to build a taxonomy and a categorized item pool. The second is an LLM-based Recommendation step, which infers user’s preference based on their historical interactions. This approach effectively handles large item pools, making it feasible to work within LLM token limits, leading to a more efficient, accurate, and scalable recommendation process. Our contributions are summarized as:
@@
Figure 2: The proposed TaxRec for zero-shot LLM-based recommendation. (a) One-time Taxonomy Categorization step aims to generate in-domain taxonomy and enrich/categorize item’s title into structured text information. (b) LLM-based Recommendation step provides ranked item lists for users based on the user’s historical interactions.
@@ 2.1 LLM for Recommendations
Recommendation systems are crucial for helping users discover relevant and personalized items Wang et al. (2024a, b); Yang et al. (2024). With the rise of LLMs, there has been growing interest in utilizing these models to improve recommendation systems Cao et al. (2024); Wang et al. (2024c). LLM-based recommenders can be broadly divided into two categories: discriminative and generative Wu et al. (2023). Discriminative methods use LLMs to learn better user and item representations from contextual information Hou et al. (2022); Li et al. (2023); Xiao et al. (2022); Zhang et al. (2022); Yuan et al. (2023); Yao et al. (2022), while generative methods leverage LLMs’ ability to generate recommendations by framing traditional ranking tasks as natural language tasks. Instead of computing scores, generative systems use techniques like prompt tuning and in-context learning to produce recommendations directly.
One of the first generative approaches was Geng et al. (2022), which utilized the pre-trained T5 model. More recent works have explored using LLMs for recommendations without fine-tuning Dai et al. (2023); Gao et al. (2023); Liu et al. (2023); Lyu et al. (2023); Wang et al. (2023c). Wang et al. (2023b) integrated databases with LLMs to act as autonomous agents for recommendation tasks, and Wang et al. (2023a) augmented user-item interaction graphs with LLMs. Lyu et al. (2023) studied the impact of different prompts on recommendation outcomes, while Wang et al. (2023c) introduced a multi-round self-reflection framework for sequential recommendations.
However, these works did not address the challenge of vague and unstructured text representations of items, which can hinder LLMs’ recommendation capabilities. In this paper, we propose a taxonomy-guided framework to address this issue.
@@ 2.2 Zero-shot Recommendations
Zero-shot recommendation has become an important area in recommendation systems, focusing on predicting user preferences without parameter adjustment. Various approaches have been developed to tackle this challenge, such as content-based methods that use item attributes Lian et al. (2018); Zhang et al. (2023b); Cao et al. (2023) and techniques that leverage pre-trained language models to extract item and user characteristics from text Ding et al. (2022); Hou et al. (2022); Li et al. (2023). Recently, LLMs have been explored for zero-shot recommendations Hou et al. (2024); Wang and Lim (2023); He et al. (2023); Feng et al. (2024); Wang et al. (2023b). However, these studies often face limitations due to the restricted context length of LLMs, making it difficult to input all items. Some approaches address this by incorporating external tools Wang et al. (2023b); Feng et al. (2024), while others use a plug-in recommendation model to narrow the candidate pool Hou et al. (2024); Wang and Lim (2023); He et al. (2023). Despite these efforts, none have fully solved this problem using LLM knowledge alone. Our work introduces a taxonomy-guided LLM recommender that compresses the item pool via an LLM-generated taxonomy, enabling effective zero-shot recommendations without external knowledge.
@@ 3.1 Problem Formulation
Without other user-item interactions, LLMs act as zero-shot recommenders when users directly seek recommendations. The task is to generate the Top-k recommended items $i$s from the candidate item pool =$\mathcal{I}=\{i_{j}\}_{j=1}^{|\mathcal{I}|}$ only based on user’s historical interactions =$\mathcal{H}=\{i_{1},i_{2},...,i_{|\mathcal{H}|}\}$ and the knowledge within LLMs. As a pure text-based approach, each item $i$ is a title string as shown in Figure 2. The task can then be represented as designing an LLM-based function:
@@ 3.2 One-time Taxonomy Categorization
The first step is a one-time generation, which aims to structure and clarify items into a categorized item pool. The original item text representation is vague and unstructured, which poses challenges for LLMs to understand and infer user’s interest. As the first item within the pool shown in Figure 2(a), “1984” can be represented as either year/book/movie. Without sufficient in-domain background knowledge, direct recommendation in zero-shot manner with these vague and unstructured textual information is challenging for LLMs.
@@
where $P_{\text{Categorization}}$ is the Categorization Prompt as shown in Table 1 to obtain $i$’s categorized feature list as $i^{C}=[f_{1},f_{2},...,f_{|i^{C}|}]$. We can structure and enrich item textual descriptions with knowledge from LLMs. For example, as shown in the categorized item pool in Figure 2(a), the book “1984” is enriched with “fiction” as genre and “power” as theme. Compared with the original vague book title, the enriched texts provide more detailed information to assist LLMs inference user’s interests. The categorized item pool $\mathcal{I}^{C}$ is obtained by categorizing items in $\mathcal{I}$ with Equation 3.
@@ 3.3 LLM-based Recommendation
where $s$ is the categorized recommendation, which is a text sequence of key-value pairs representing item’s features within $\mathcal{T}$. $P_{\text{Recommendation}}$ is the recommendation prompt given $\mathcal{H^{C}}$ and $\mathcal{T}$as shown in Table 1. Using $\mathcal{T}$ instead of the item pool can greatly decrease the prompt length and fit the in-domain item’s information within the limited context requirement from LLMs. $P_{\text{Recommendation}}$ also regularizes LLM’s generation format as a list of features based on $\mathcal{T}$. $s$ is further parsed as the feature list $F=[f_{1},f_{2},...,f_{|F|}]$ representing recommended features. Then the ranking score of each item $i$ is calculated as:
@@
Then items with Top-k highest ranking scores are retrieved from the item pool and recommended to users. In summary, we designed a framework TaxRec, which uses a taxonomy as the intermediate, to unify the representation of items throughout the recommendation pipeline. TaxRec can retrieve LLM’s knowledge for zero-shot recommendations without any training and other users’ interactions with item.
@@ 4.1.1 Datasets
Datasets Methods R@1 R@5 R@10 N@1 N@5 N@10
Movie Popularity 0.005 0.035 0.160 0.005 0.020 0.061
AvgEmb 0.000 0.040 0.100 0.000 0.020 0.039
ZESRec 0.032 0.095 0.222 0.032 0.059 0.099
UniSRec 0.032 0.063 0.143 0.032 0.048 0.074
RecFormer 0.016 0.141 0.219 0.016 0.077 0.103
DirectRec-LLaMA2 0.033 0.058 0.085 0.033 0.042 0.051
TaxRec-LLaMA2 0.045 0.126 0.190 0.045 0.095 0.148
DirectRec-GPT4 0.045 0.100 0.180 0.045 0.074 0.099
TaxRec-GPT4 0.060 0.175 0.300 0.060 0.117 0.157
Book Popularity 0.030 0.070 0.155 0.030 0.046 0.073
AvgEmb 0.005 0.075 0.115 0.005 0.038 0.051
ZESRec 0.005 0.070 0.115 0.005 0.037 0.051
UniSRec 0.000 0.050 0.085 0.000 0.025 0.035
RecFormer 0.010 0.060 0.125 0.010 0.033 0.054
DirectRec-LLaMA2 0.001 0.010 0.015 0.001 0.004 0.006
TaxRec-LLaMA2 0.040 0.099 0.150 0.040 0.072 0.109
DirectRec-GPT4 0.000 0.015 0.025 0.000 0.006 0.010
TaxRec-GPT4 0.070 0.150 0.240 0.070 0.109 0.138
Table 2: Performance comparison between different zero-shot recommendation baselines and TaxRec. We report Recall(R) and NDCG(N) @(1, 5, 10) results multiplied by 10. The boldface indicates the best result and the underlined indicates the second best. All TaxRec results are significantly better than the baselines with $p<0.05$.